OpenStudy (ashleyy_smith):

Need help with 2-3 questions please (: , Two less than the product of 5 and a number is equal to 6. ~~Use the variable y for the unknown number~~ Have to turn it into an equation..

OpenStudy (jiteshmeghwal9):

Product of number y & 5=5y if we subtract 2 from 5y, we get 6. can u form it now ?

OpenStudy (ashleyy_smith):

hmm.. 2 - 5y =6?

OpenStudy (jiteshmeghwal9):

we need to subtract 2 from 5y not 5y from 2

OpenStudy (ashleyy_smith):

so I put the y with the 2?

OpenStudy (jiteshmeghwal9):

If u want to subtract 2 from 5 u write 5-2 & if u want to subtract 5 from 2 u write 2-5. Do u get it now wht i m telling u ?

OpenStudy (nincompoop):

Two less than the product of 5 and a number is equal to 6. Use the variable y for the unknown number -2 + (5 * y) = 6 -2 + 5y = 6 it does not matter in which order you assemble the worded problem into expression so long as you follow the signs properly.

OpenStudy (ashleyy_smith):

OHHH alright. You worded it easier for me.

OpenStudy (ashleyy_smith):

Next one is, Twice the sum of a number and 4 is equal to 5. 2b+4=5? right?

OpenStudy (nincompoop):

In the problem, the bottom line is that 2 is being subtracted into the product of 5 and y (a number). It will be typically written as 5y - 2 = 6 but that is the same as -2 + 5y = 6 look at this: +5y - 2 = 6 then this -2 + 5y = 6 if the leading (first) coefficient of the expression is positive, that is with a plus sign, we omit the + by convention. Therefore 5y -2 = 6 is how it is usually written.

OpenStudy (nincompoop):

TWICE the sum so we are multiplying 2 by some sum 2 * ( ____ + ____ ) = Twice the sum of a number and 4 is equal to 5 2 * ( ____ + 4 ) = 5 Then replace the blank by a variable

OpenStudy (nincompoop):

This is going to take a lot of practice so do not be afraid if you are making mistakes. It is part of the learning process to make mistakes.

OpenStudy (ashleyy_smith):

2(b+4)=5 right?

OpenStudy (nincompoop):

ye

OpenStudy (ashleyy_smith):

And thank you for understanding (:

OpenStudy (ashleyy_smith):

The difference of a number divided by 2 and 6 equals 8.. I know its subtraction..

OpenStudy (ashleyy_smith):

@nincompoop

OpenStudy (nincompoop):

try

OpenStudy (nincompoop):

even for me that is a little tricky

OpenStudy (ashleyy_smith):

right?!?

OpenStudy (nincompoop):

well there is a subtraction somewhere, but there's also division

OpenStudy (ashleyy_smith):

can it be -w (thats the variable)

OpenStudy (nincompoop):

okay